OpenStudy (anonymous):

SOLVE THE EQUATION 8.3-y=-5.6

OpenStudy (anonymous):

Did you attempt it?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Do u change -y=-5.6 into a positive number?

OpenStudy (anonymous):

What happened to the 8.3?

OpenStudy (anonymous):

Sry wrong question

OpenStudy (anonymous):

hrm?

OpenStudy (anonymous):

what is the answer to 8.3-y=2.7

OpenStudy (anonymous):

What have you done so far?

OpenStudy (anonymous):

What does "=" mean?

OpenStudy (anonymous):

i have the answer of -y=-5.6

OpenStudy (anonymous):

do u change it into a positive number

OpenStudy (anonymous):

Well, why don't you try adding y + 5.6 to both sides.

OpenStudy (anonymous):

And see what happens.

OpenStudy (anonymous):

Or you can multiply both sides by -1 if you are good with multiplying by negatives.

OpenStudy (anonymous):

ok its y=5.6

OpenStudy (anonymous):

Yeah.

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

8.3-y=-5.6 -8.3 -8.3 y=5.6

OpenStudy (anonymous):

that how you slove it